OpenStudy (anonymous):

I need help with a long Radical equation all i need to simplify it all the way down, can anyone help me?

OpenStudy (anonymous):

\[9 \sqrt[4]{10} + 2 \sqrt[3]{10} -6 \sqrt[4]{10} -6 \sqrt[3]{10}\]

OpenStudy (zehanz):

This expression has no common factors. What we can do, is add/subtract like terms: there are two terms with \(\sqrt[4]{10}\) and two terms with \(\sqrt[3]{10}\), so they can be written as one... Begin with writing it as: \((9\sqrt[4]{10}-6\sqrt[4]{10})+(2\sqrt[3]{10}-6\sqrt[3]{10})\), now simplify.

OpenStudy (anonymous):

Oh, so now foil?

OpenStudy (zehanz):

No, there is nothing to FOIL, the brackets are there to help you see the two different types of radicals. What is \(9\sqrt[4]{10}-6\sqrt[4]{10}\) ?

OpenStudy (anonymous):

\[3\sqrt[4]{10}\]

OpenStudy (zehanz):

Right!. Now the other; what will that be?

OpenStudy (anonymous):

\[-4\sqrt[3]{10}\]

OpenStudy (zehanz):

Right again! So you have these two amounts of different kind of radicals: \(3\sqrt[4]{10}-4\sqrt[3]{10}\). So that's it! You're done!

OpenStudy (anonymous):

AWesome THanks mane your a huge help!

OpenStudy (zehanz):

YW!